A straight wire of length 0.75 m carries a current of 35 A. The wire is at right angles to a magnetic field of strength 0.058 T. Calculate the force on the wire.


1
Expert's answer
2021-12-28T10:59:18-0500

Answer

Given data as below

Length of wire L=0.75m

Current in wire I=35A

magnetic field of strength B=0.058 T

Right angle "\theta=90°"

Force on this wire given by

"F=ILB \sin\theta"

Putting all values

"F=35\times0.75\times0.058\times\sin90°"

"F=26.25\times0.058\times1\\F=1.5225N"

So force on this wire is 1.5225N.